OpenStudy (anonymous):

Solve ln(2x-1)=8 round to nearest thousandth a.1,489.979 b.2,979.958 c.2,981.458 d.1,490.979

OpenStudy (anonymous):

Raise both sides by the base e: 2x - 1 = e^8 2x = e^8 - 1 x = (e^8 - 1)/2 which is rounded to 1,489.979

OpenStudy (anonymous):

What does the e mean?

OpenStudy (anonymous):

e is the Euler's number, 2.7...

OpenStudy (anonymous):

o ok